OpenStudy (anonymous):

another part of the project need steps to x^-7

OpenStudy (alexandervonhumboldt2):

look: a^(-b)=1/(a^b)

OpenStudy (alexandervonhumboldt2):

so hmm, is there any given value for x?

OpenStudy (anonymous):

all it says is x^ of negative 7

OpenStudy (alexandervonhumboldt2):

hm well i think they just mean that x^-7=1/(x^7)

OpenStudy (anonymous):

can you do the steps to that we need the steps :)

OpenStudy (alexandervonhumboldt2):

this is the rule: a^-b=1/a^b. it is the rule, there are no steps.

OpenStudy (alexandervonhumboldt2):

instead of a there is x and instedad of b there is 7

OpenStudy (anonymous):

Thankyou so much for your help :)

OpenStudy (alexandervonhumboldt2):

np